# Goal-Oriented Communications for Real-time Inference with Two-Way Delay

### v1

@@ Abstract
We design a goal-oriented communication strategy for remote inference, where an intelligent model (e.g., a pre-trained neural network) at the receiver side predicts the real-time value of a target signal based on data packets transmitted from a remote location. The inference error depends on both the Age of Information (AoI) and the length of the data packets. Previous formulations of this problem either assumed IID transmission delays with immediate feedback or focused only on monotonic relations where inference performance degrades as the input data ages. In contrast, we consider a possibly non-monotonic relationship between the inference error and AoI. We show how to minimize the expected time-average inference error under two-way delay, where the delay process can have memory. Simulation results highlight the significant benefits of adopting such a goal-oriented communication strategy for remote inference, especially under highly variable delay scenarios.
@@ I Introduction
Intelligent models, e.g. neural networks, have become powerful tools for solving complex computational problems, thanks to advancements in artificial intelligence. These models have not only increased the popularity of today’s most up-to-date technologies but have also played a key role in enabling their current potential. Digital twins, industrial robots, and self-driving cars are prominent examples of such technologies.
An important application of digital twins over long-distance network connections is the monitoring of spacecraft. Consider training a digital twin on Earth based on data coming from rovers on Mars or calibrating the virtual replica of the spacecraft/satellite using live telemetry data [2, 3]. Digital twins have been used in spacecraft design and operations, ranging from assisting ground control in saving the Apollo 13 astronauts to testing thermal protection on the Space Shuttle.
@@
In recent years, digital twins have also gained importance in the monitoring of processes in industrial environments, especially in hazardous or hard-to-reach environments or when it is impractical to be on site for monitoring. One example is the digital twin implementation for the METU high voltage testing laboratory. [^1] A visual of this implementation, showing a screenshot of the digital twin of the METU high voltage testing laboratory on the left, and a photograph of the actual site on the right, can be accessed at https://cng-eee.metu.edu.tr/wp-content/Digital_twin.pdf?_t=1726991660. The digital twin can be monitored over the Internet, and device settings in the physical experiment configuration on site can be remotely changed on demand directly on the twin.
@@
Such systems, involving real-time monitoring, training, etc., often require computation that relies on data sent from remote locations to maintain good performance. Motivated by this, we focus on goal-oriented communication design for the remote inference problem [4, 5]. Goal-oriented communication has been recently introduced to address the scalability problem in next-generation communication networks [6]. This concept emphasizes that networks should focus on the effective accomplishment of the task at the destination rather than merely addressing the transmission problem [7], which aims to reliably transmit data produced by a source. Solving the effective communication problem with efficient use of network resources requires combining the data generation and transmission processes, ensuring that data samples most significant to the computation at the destination are delivered in a timely manner [8]. This implies that the communication link should select which samples to transmit based on the state of the network (e.g., the delay state). Recent efforts in the communication and control communities highlight the value of this approach in reducing communication requirements while maintaining application performance [9, 10, 11, 12, 13, 14].
@@
Since we cannot expect the link layer to operate jointly with the application layer for all possible applications, a surrogate metric that the link layer can utilize when deciding which samples to select and transmit is useful. The Age of Information (AoI) at time $t$, denoted by $\Delta(t)$, is defined in [15] as $\Delta(t)=t-U_{t}$, where $U_{t}$ is the generation time of the most recently delivered data packet. We utilize the AoI as an intermediate metric that can be monitored by the link layer, independent of the specific application. A mapping of application performance in relation to the AoI is all that is required for the link layer to function in a goal-oriented manner. This work builds upon the findings presented in the series of papers [4, 16, 5, 17, 1], which first demonstrated the effectiveness of the AoI metric as a surrogate for enhancing communication in the remote tracking or inference of a process.
In this paper, we study a remote inference problem, where an intelligent model infers the real-time value of a target signal based on data transmitted from a remote location. We assume that the transmitter regularly samples a source and forms data packets containing a number of data samples, which we refer to as the packet length, before each transmission. The intelligent model on the receiver side performs inference using these data packets. The inference error is a function of the AoI and the length of the data packet used by the intelligent model. Recent studies [4, 5] have shown that, in various remote inference problems, the inference error (4) does not always increase monotonically with AoI. This contradicts the conventional assumption that a packet with AoI $=0$ necessarily yields better performance than one with AoI $>0$. Therefore, rather than the generate-at-will model [18, 19], we adopt the “selection-from-buffer” model proposed in [4, 5], which enables the scheduler to choose either fresh or stale data samples from the buffer. Moreover, the inference error decreases monotonically as the packet length increases [17]. This improvement comes at the cost of higher resource usage and increased delay. In our communication design, we explore both the possibly non-monotonic relationship between inference error and AoI, as well as the interplay between packet length and transmission delay.
@@
Due to the need for extreme connectivity, next-generation communication networks are expected to involve numerous connections, leading to the existence of multiple routes between any nodes in the network. The relays in such networks may enforce specific routes for certain data flows, which can dynamically change over time to meet the varying demands of multiple flows. For instance, consider a network involving both terrestrial and non-terrestrial connections. A flow can be fully serviced by non-terrestrial connections, terrestrial connections, or a combination of both, depending on conditions such as the priority of the flow, network congestion, or the availability of satellites. Another example is busy industrial environments, where communication routes may be intermittently unavailable due to shadowing from moving objects, causing flows to be rerouted over time. Consequently, goal-oriented communication design for remote inference must be adaptable to delay conditions significantly varying with memory.
@@
  • Extending the system models in [5, 17], we formulate and solve a learning and communication co-design problem for remote inference under two-way delay that varies significantly with memory. We consider both time-invariant and time-variable packet length selection scenarios to address the varying computational capabilities of practical systems. The derived optimal scheduling policies minimize the time-average inference error for a given, possibly non-monotonic, inference error function corresponding to a particular remote inference system, thereby making the communication goal-oriented.
  • We model the learning and communication co-design problem under time-invariant packet length selection as an infinite-horizon average-cost semi-Markov decision process (SMDP). Such problems are often solved using dynamic programming [20, 21] and typically do not have closed-form solutions. However, we derive a closed-form solution for this scenario, presented in Theorem 1.
@@
  • We conduct two experiments to evaluate the performance of the derived optimal scheduling policies: (i) remote inference of autoregressive (AR) processes, which provides a model-based evaluation; and (ii) cart-pole state prediction, which offers a trace-driven evaluation. Simulation results indicate the substantial benefits of adopting a goal-oriented communication strategy for remote inference.
@@ I-A Related Work
The works most closely related to this paper are [5, 17], which developed scheduling policies for remote inference, considering a possibly non-monotonic dependency between AoI and practical performance. While [5] developed scheduling policies for fixed packet length, the paper [17] jointly optimized packet length and scheduling strategies. Both the paper [5, 17] considered random IID delay for packet transmissions from the transmitter to the receiver and assumed delay-free feedback from the receiver to the transmitter. In practical systems, a non-zero feedback delay exists [40, 41], which affects the performance of the remote inference algorithm. Moreover, due to the envisioned structure of next-generation communication networks, which contain multiple routes between any pair of nodes, the IID delay may not be a good model. This necessitates a model that captures the memory in the delay process, as in [19]. Extending the problem formulations in [5, 17], this paper designs a goal-oriented communication strategy that minimizes the average inference error for remote inference under two-way delay significantly varying with memory.
@@ II System Model
We consider a remote inference system, which comprises a data source, a transmitter, and a receiver, as illustrated in Fig. 1. On the transmitter side, the source signal $V_{t}$ is regularly sampled from the data source, and the samples are stored in a buffer containing the most recent $B$ samples. The scheduler determines when to transmit and which samples in the buffer to transmit. On the receiver side, a predictor, e.g., a pre-trained neural network, infers the real-time value of a target signal $Y_{t}$ based on the most recently received samples. Such remote inference systems are crucial for many real-time applications, including sensor networks, airplane/vehicular control, robotics networks, and Cyber-Physical Systems (CPS), where timely and accurate remote monitoring and control are essential for maintaining system stability and performance.
Figure 1: A remote inference system, where the packet $X_{t-b}^{l}$ represents the group of samples $\big{(}V_{t-b},V_{t-b-1},\ldots,V_{t-b-l+1}\big{)}$.
We assume that the system is time-slotted, and it starts operation at time slot $t=0$. We adopt the selection-from-buffer model proposed in [5], where, at each time slot $t$, a new sample of the source signal $V_{t}\in\mathcal{V}$ is added to the buffer while the oldest sample is discarded. The buffer, with size $B$, is assumed to be initially full, storing the samples $\big{(}V_{0},V_{-1},\ldots,V_{-B+1}\big{)}$. Thus, at any time slot $t$, the buffer contains the most recent $B$ samples, $\big{(}V_{t},V_{t-1},\ldots,V_{t-B+1}\big{)}$. Before the $i$-th transmission, the scheduler selects a group of consecutive samples from the buffer and forms the packet $X_{t-b_{i}}^{l_{i}}=\big{(}V_{t-b_{i}},V_{t-b_{i}-1},\ldots,V_{t-b_{i}-l_{i}+1}\big{)}\in\mathcal{V}^{l}$. Here, $l_{i}\in\{1,\ldots,B\}$ and $b_{i}\in\{0,\ldots,B-l_{i}\}$ represents the packet length and buffer position of the $i$-th packet $X_{t-b_{i}}^{l_{i}}$, respectively. Fig. 2 indicates a sample evolution for two consecutive transmission periods. The packet $X_{t-b_{i}}^{l_{i}}$ is submitted at time slot $S_{i}$ and delivered to the receiver at time slot $D_{i}$. Upon delivery, the receiver sends an acknowledgment (ACK) back to the transmitter, which is received at time slot $A_{i}$. We assume that the transmitter always waits for feedback before submitting a new packet. In other words, the transmitter remains silent during the time slots between $S_{i}$ and $A_{i}$ for all $i$.
@@ II-A Communication Network Model
The remote inference system operates on top of a communication network with multiple routes $\{1,2,\ldots,N\}$ connecting the transmitter and the receiver. Each packet submitted into the network is directed to a route based on the network status regarding availability and congestion. If a packet is transmitted through a specific route, its corresponding feedback is also transmitted through that same route. Each route has distinct transmission and feedback delay distributions, and the transmission delay also varies with the packet length. The delay is at least one time slot for each packet or ACK transmission.
Let the $i$-th epoch consist of the time slots in the interval $[A_{i-1},A_{i})$. Define $C_{i}$ as the delay state in the $i$-th epoch, with $c_{i}\in\{1,2,\ldots,N\}$ denoting its realization. The delay state $C_{i}$ represents the network route used in the $i$-th epoch and specifies the transmission and feedback delay distributions. We assume that $C_{i}$ evolves according to a finite-state ergodic Markov chain with transition probabilities $p_{ij}$, where $i,j\in\{1,2,\ldots,N\}$, and $p_{ij}$ represents the probability of transitioning from state $i$ to state $j$. The Markov chain makes a single transition at time slot $A_{i}$ and none otherwise. Let $T_{i}(l_{i})=D_{i}-S_{i}$ and $F_{i}=A_{i}-D_{i}$ represent the transmission and feedback delays incurred in the $i$-th epoch, respectively.
The receiver detects the delay state $c_{i}$ at each delivery time $D_{i}$ and embeds this information into the ACK message. The ACK notifies the scheduler that the previous packet transmission is complete and that the next packet can be submitted to the network. The network is reliable, meaning no packet or message is lost during transmission.
II-B Inference Error
The Age of Information (AoI) on the receiver side, $\Delta(t)$, is the time difference between the current time $t$ and the generation time $t-\Delta(t)$ of the freshest sample $V_{t-\Delta(t)}$ in the most recently delivered packet $X_{t-\Delta(t)}^{l}$. The AoI at time slot $t$ is determined by

### v3

# Goal-Oriented Status Updating for Real-time Remote Inference over Networks with Two-Way Delay

@@ Abstract
We study a setting where an intelligent model (e.g., a pre-trained neural network) predicts the real-time value of a target signal using data samples transmitted from a remote source according to a scheduling policy. The scheduler decides on i) the age of the samples to be sent, ii) when to send them, and iii) the length of each packet (i.e., the number of samples contained in each packet). The dependence of inference quality on the Age of Information (AoI) for a given packet length is modeled by a general relationship. Previous work assumed i.i.d. transmission delays with immediate feedback or were restricted to the case where inference performance degrades as the input data ages. Our formulation, in addition to capturing non-monotone age dependence, also covers Markovian delay on both forward and feedback links. We model this as an infinite-horizon average-cost Semi-Markov Decision Process. We obtain a closed-form solution that decides on (i) and (ii) for any constant packet length. The solution for when to send is an index-based threshold policy, where the index function is expressed in terms of the delay state and AoI at the receiver. The age of the packet selected is a function of the delay state. We separately optimize the value of the constant length. We also develop an index-based threshold policy for the variable length case, which allows a complexity reduction. In simulation results, we observe that our goal-oriented scheduler drops inference error down to one sixth with respect to age-based scheduling of unit-length packets.
@@ I Introduction
Today, many previously complex computational problems enjoy the aid of AI. Among these AI-aided approaches are intelligent models, e.g., neural networks. These models have played a key role in enabling certain technologies as digital twins, industrial robotics, and self-driving cars. However, these models rely on data to be delivered to them, which in the remote computation context, depends on the network delivering the relevant data in a timely fashion. This is especially interesting in the case of unavoidable and time-varying transmission delays and is within the realm of Goal-oriented Networking for AI [2].
@@
An extreme example of monitoring over long-distance network connections is a spacecraft digital twin. Consider training a digital twin on Earth based on data coming from rovers on Mars or calibrating the virtual replica of the spacecraft/satellite using live telemetry data [3, 4]. Furthermore, in recent years, digital twins have also gained importance in the monitoring of processes in industrial environments, especially in hazardous or hard-to-reach environments or when it is impractical to be on site for monitoring. [^1] A screenshot of the digital twin of the METU high voltage testing laboratory is available at https://cng-eee.metu.edu.tr/wp-content/Digital_twin.pdf?_t=1726991660. The digital twin can be monitored over the Internet, and device settings in the physical experiment configuration on site can be remotely changed on demand directly on the twin.
@@
We focus on goal-oriented communication design for real-time estimation and inference performed by intelligent models [5, 6], which are often essential for such applications. Goal-oriented communication has been recently introduced to address the scalability problem in next-generation communication networks. This concept emphasizes that networks can facilitate the effective accomplishment of the task at the destination rather than merely addressing the transmission problem [7], which aims to reliably transmit data produced by a source. Solving the effective communication problem with efficient use of network resources requires combining the data generation and transmission processes, ensuring that data samples most significant to the computation at the destination are delivered in a timely manner [8]. This implies that the communication link should select which samples to transmit based on the state of the network (e.g., the delay state). Recent efforts in the communication and control communities highlight the value of this approach in reducing communication requirements while maintaining application performance [9, 10, 11, 12, 13, 14].
@@
Age of Information (AoI) at time $t$, denoted by $\Delta(t)$, is defined in [15] as $\Delta(t)=t-U_{t}$, where $U_{t}$ is the generation time of the most recently delivered data packet. The application-layer real-time estimation and inference performance is mapped as a function of the AoI. This function is all that is required for the link layer to operate in a goal-oriented manner. This work builds upon the findings presented in a series of papers [5, 16, 6, 17, 1, 18, 19, 20], which first demonstrated the effectiveness of AoI function as a surrogate metric for enhancing timely communication for remote estimation and inference.
In this paper, we study a remote inference problem, where an intelligent, pre-trained model at the receiver is utilized to infer the real-time value of a target signal, $Y_{t}$, based on a data packet, $X_{t-\delta}^{l}=(V_{t-\delta},V_{t-\delta-1},\ldots,V_{t-\delta-l+1})$, transmitted from a remote location. The data packet $X_{t-\delta}^{l}$ comprises samples of the source signal $V_{t}$ at the transmitter. Here, $\delta$ represents the AoI of the freshest sample in the packet, and $l$ denotes the packet length, corresponding to the number of samples included in the packet.
The inference error is characterized as a function of both the AoI and the packet length. Recent studies [5, 6] have demonstrated that, for various remote inference systems, the relationship between the inference error and AoI for a given packet length can be non-monotonic. This finding contradicts the conventional assumption that a packet with ${\text{AoI}=0}$ necessarily yields better performance than one with ${\text{AoI}>0}$. Consequently, rather than adopting the generate-at-will model [21, 18], we use the “selection-from-buffer” model proposed in [5, 6], which enables the scheduler to choose either fresh or stale data samples from the buffer.
Such non-monotonicity between the inference error and AoI arises when the source signal, $V_{t}$, and the target signal, $Y_{t}$, exhibit a delayed relationship, such as $Y_{t}=f(V_{t-\Upsilon})$ [6, 22, 20]. For instance, $\Upsilon$ may represent a combination of the communication delay between the controller and the actuator and the actuation delay in a physical networked control system. An example of this scenario is illustrated in the leader-follower robot setup in [6, Fig. 2]. Additionally, non-monotonicity also occurs when the relationship between $V_{t}$ and $Y_{t}$ exhibits periodicity, as shown in the temperature prediction example in [6, Fig. 4]. In such cases, when communication delays prevent the transmission of sufficiently fresh samples to the receiver, transmitting aged samples can exploit the periodicity to enhance performance.
Moreover, the inference error is non-increasing with packet length for a given AoI [17], as additional information can only improve the estimation accuracy. However, this improvement comes at the cost of longer transmission delays. In our communication design, we investigate the interplay between packet length and transmission delay, as well as the possibly non-monotonic relationship between inference error and AoI.
@@
Finally, next-generation communication networks are expected to accommodate numerous connections, resulting in multiple routes between any pair of nodes. The relays in such networks may enforce specific routes for certain data flows, which can dynamically change over time to address the varying demands of multiple flows. For instance, consider a network involving both terrestrial and non-terrestrial connections. A flow may be serviced entirely by non-terrestrial connections, terrestrial connections, or a combination of both, depending on factors such as the flow’s priority, network congestion, or satellite availability. Therefore, goal-oriented communication design must be adaptable to delay conditions that vary significantly with memory.
@@
  • We extend the system models in [6, 17] to more practical scenarios by considering Markovian delay on both forward and feedback links. Based on this extension, we formulate and solve a learning and communication co-design problem for real-time remote inference. We consider both time-invariant and time-variable packet length selection to address the varying computational capabilities of practical systems. The derived optimal scheduling policies minimize the time-average inference error for a given, possibly non-monotonic, inference error function corresponding to a particular remote inference application, thereby making the communication goal-oriented.
  • We model the learning and communication co-design problem under time-invariant packet length selection as an infinite-horizon average-cost Semi-Markov Decision Process (SMDP). Such problems are often solved using dynamic programming [23, 24] and typically do not have closed-form solutions. However, we derive a closed-form solution to this problem for any constant packet length, presented in Theorem 1. The solution consists of two main parts: (i) the AoI of data packets to be sent are determined as a stationary function of the network state, i.e., the delay state, and (ii) the transmission instants are governed by an index-based threshold policy. The index function is expressed in terms of the delay state of the network and the AoI at the receiver. We then separately optimize the value of constant packet length.
@@
  • We conduct two experiments to evaluate the performance of our optimal scheduling policies: (i) remote inference of auto-regressive (AR) processes, offering a model-based evaluation; and (ii) cart-pole state prediction, providing a trace-driven evaluation. In simulation results, we observe that our goal-oriented scheduler drops inference error down to one sixth with respect to age-based scheduling of unit-length packets.
@@ I-A Related Work
The works most closely related to this paper are [6, 17], which developed scheduling policies for remote inference, considering a possibly non-monotonic dependency between AoI and practical performance. While [6] developed scheduling policies for fixed packet length, [17] jointly optimized packet length and scheduling strategies. However, both studies [6, 17] assumed random i.i.d. delay for packet transmissions and immediate feedback. This paper extends the problem formulations in [6, 17] to more practical scenarios and develops a goal-oriented communication strategy that minimizes the average inference error for remote inference under two-way delay (i.e., incorporating random feedback delay [40, 41]) that varies significantly with memory.
@@ II System Model
Figure 1: A remote inference system that adopts the selection-from-buffer medium access model [6, 5].
We consider a time-slotted remote inference system, which comprises a data source, a transmitter, and a receiver, as illustrated in Fig. 1. On the transmitter side, the source signal $V_{t}$ is regularly sampled in each time slot $t$. The samples are stored in a buffer containing the most recent $B$ samples $\big{(}V_{t},V_{t-1},\ldots,V_{t-B+1}\big{)}$, waiting to be sent to the receiver. On the receiver side, a predictor, e.g., a pre-trained neural network, is used to infer the real-time value of a target signal $Y_{t}$ of interest. In practical time-series forecasting algorithms, the predictor needs to use multiple consecutive samples, e.g., $X_{t-b}^{l}=(V_{t-b},V_{t-b-1},\ldots,V_{t-b-l+1})$, to infer $Y_{t}$. We will design a transmission scheduler that determines (i) when to send a group of consecutive samples over the communication network to the receiver and (ii) which group of consecutive samples in the buffer to send. This medium access model is called “selection-from-buffer,” which was proposed recently in [5, 6]. Such remote inference systems are crucial for numerous real-time applications, including sensor networks, airplane/vehicular control, robotics networks, and digital twins where timely and accurate remote monitoring and control are essential for maintaining system stability and performance.
We assume that the system starts operation at time slot $t=0$. In the selection-from-buffer model, a new sample of the source signal $V_{t}\in\mathcal{V}$ is added to the buffer at the beginning of each time slot $t$, and the oldest sample is discarded. The buffer, with size $B$, is assumed to be initially full, storing the samples $\big{(}V_{0},V_{-1},\ldots,V_{-B+1}\big{)}$. Thus, at any time slot $t$, the buffer contains the most recent $B$ samples $\big{(}V_{t},V_{t-1},\ldots,V_{t-B+1}\big{)}$. The transmitter sends status update packets to the receiver one by one. Let $S_{i}$ and $D_{i}$ represent the transmission start time and delivery time of the $i$-th transmitted packet, respectively, such that $S_{i}<D_{i}$. At time slot $t=S_{i}$, the scheduler selects a group of consecutive samples from the buffer and forms the packet $X_{S_{i}-b_{i}}^{l_{i}}=\big{(}V_{S_{i}-b_{i}},V_{S_{i}-b_{i}-1},\ldots,V_{S_{i}-b_{i}-l_{i}+1}\big{)}\in\mathcal{V}^{l_{i}}$. Here, $l_{i}\in\{1,\ldots,B\}$ represents the sequence length of the $i$-th packet $X_{S_{i}-b_{i}}^{l_{i}}$ while $b_{i}$ represents the relative position of the sample $V_{S_{i}-b_{i}}$ in the buffer. When $B=1$, the selection-from-buffer model reduces to the generative-at-will model which has been studied extensively in previous works [37, 18, 38, 39, 19, 27, 28, 29]. Hence, the packet $X_{S_{i}-b_{i}}^{l_{i}}$ is submitted to the communication network at time slot $S_{i}$ and delivered to the receiver at time slot $D_{i}$. Upon delivery, the receiver sends an acknowledgment (ACK) back to the transmitter, which is received at time slot $A_{i}$. We assume that the transmitter always waits for feedback before submitting a new packet. In other words, the transmitter remains silent during the time slots between $D_{i}$ and $A_{i}$ such that $S_{i}<D_{i}<A_{i}\leq S_{i+1}$ for all $i$. Fig. 2 provides an illustration of the $i$-th and the $(i+1)$-th packet transmission epochs.
@@ II-A Communication Network Model
Figure 2: An illustration of the $i$-th and the $(i+1)$-th packet transmission epochs.
Consider that the communication network may have multiple routes $\{1,2,\ldots,N\}$ connecting the transmitter and the receiver. The network directs each submitted packet to one of these available routes. If a packet is transmitted through a specific route, its corresponding feedback is also transmitted through that same route.
Let the $i$-th packet transmission epoch consist of the time slots $A_{i-1},A_{i-1}+1,\ldots,A_{i}-1$. Define $T_{i}(l_{i})=D_{i}-S_{i}$ as the transmission delay incurred in the $i$-th epoch and $F_{i}=A_{i}-D_{i}$ as the feedback delay. Additionally, let $C_{i}$ represent the delay state in the $i$-th epoch, with $c_{i}\in\{1,2,\ldots,N\}$ denoting its realization. The delay state $C_{i}$ is determined by the route used in the $i$-th packet transmission epoch. Each route has distinct transmission and feedback delay distributions, and the transmission delay also varies with the packet length. The delay is at least one time slot for each packet or ACK transmission. We assume that $C_{i}$ evolves according to a finite-state ergodic Markov chain with transition probabilities $p_{ij}$, where $i,j\in\{1,2,\ldots,N\}$, and $p_{ij}$ represents the probability of transitioning from state $i$ to state $j$. The Markov chain makes a single transition at time slot $A_{i}$ and none otherwise.
The receiver detects the delay state $C_{i}$ at the packet delivery time $D_{i}$ and embeds this information into the ACK message. The ACK notifies the transmitter of (i) the delay state experienced during the previous packet transmission, (ii) the completion of the previous packet transmission, and (iii) that the next packet can be submitted to the network. The network is reliable, meaning no packet is lost during transmission.
II-B Inference Error as a Function of AoI and Packet Length
Age of Information (AoI) on the receiver side, $\Delta(t)$, is the time difference between the current time $t$ and the generation time $t-\Delta(t)$ of the freshest sample $V_{t-\Delta(t)}$ in the most recently delivered packet $X_{t-\Delta(t)}^{l}$ [15]. The AoI at time slot $t$ is determined by